Operations Report – February 08th

Crew 221 Operations Report 07Feb2022

SOL: 6

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: off (at 8pm) waiting for SOC to reach 80% to be turned on

Hours run: 11:00

From what time last night: 8:30pm

To what time this morning: 7:30am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 89%

Diesel Reading – 97%

Station Propane Reading – 60%

Ethanol Free Gasoline – gallons

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 285 gals

Static to Loft Pump used – yes

Water in GreenHab – 227.3 gals

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: (before EVA) : 129.9

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging) 75%

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: yes

Hours: 136.1

Beginning charge: 100%

Ending charge: 72%

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: operational, lose cable on suit 7 fixed again but it might neeed a more permanent fix for the lose cable. Suit 10 is not charging, voltage at battery reads 5V, battery change is recomended and permission to perfome the replacement is requested.

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment continues, now watered with 2.5l/day. watering of the plants carried out as per procedure

Summary of ScienceDome operations: 2 incubators runing, microbiology experiments going on and cultures are bieng grown in agar plates. Detection of K, P and N carried out for tree soil samples currenlty being used at the greenhab for the black corn experiment.

Summary of RAM operations: robot assembly is ongoing adn expected to continue, antenna asembly for radio comunication experiment took place today and is expected to continue.

Summary of any observatory issues: nominal

Summary of health and safety issues: strong gas smell reported and gas detector brough into de the hab, it was determined that the gas came from the oven which was not working properly.

Questions, concerns and requests to Mission Support: the oven was reconfigured to work with propane as the previous confifguration was for natural gas, and that was deteremined to be the cuase of the smell reported yesterday.

Greenhab Report – February 08th

Crew 221 GreenHab Report 08FEB2020

GreenHab Officer: Jessica Todd

Environmental control: Heating. GreenHab door shade cloth (40%) (30%) on. Door opened today for four hours from 2:45 pm to 6:52 pm due to very warm weather.

Average temperatures: 24.33°C

Hours of supplemental light: 19:30 to 00:00

Daily water usage for crops: 14.7 gallons

Daily water usage for research and/or other purposes: 1 gallon

Water in Blue Tank: 227.3 gallons

Time(s) of watering for crops: 2

Changes to crops: One of our succulents has died (picture attached). I managed to rescue three out of five, waiting on one, but sadly the other one is completely dead. Rest of the GreenHab looks really healthy, including the tomatoes. Ginger plant going strong and we have some new bell peppers and lettuce sprouting!

Narrative: Morning watering was a little late today due to the early EVA, but plants all look really healthy. Commander Maria and I set up a new experiment in the GreenHab, attempting to grow mungbeans in different ratios of Earth and Martian soil. We have several beds of mungbeans growing in 100% Earth Soil, 100% Martian Soil, 50%/50% Martian/Earth Soil and 70%/30% Martian/Earth soil. We will be monitoring the soil pH and nutrients to see if the Earth soil can populate the Martian soil with some healthy bacteria (Mark Watney style!). Evening watering was uneventful, all the plants save the succulent look pretty healthy. We should be ok to harvest herbs again tomorrow.

Harvest: NA

Support/supplies needed: NA

Other Notes: Sprayed one of the lettuce plants with pesticide. The bugs seemed to have fled the sprayed herbs and onto the lettuce. The lettuce plants in the hanging Ikea pots are having trouble with over-moist soil, as the pots have no drainage. Am I able to mix some perlite in with them to help absorb up the water? Also should I fertilize the plants at all while I’m here?


EVA Report – February 08th

Crew 221 EVA Report 08FEB-2020

EVA # 7

Author: Jessica Todd (GreenHab Officer)

Purpose of EVA: The primary purpose was to determine test sites for our Mars rover. EVA#6 photographed some interesting areas of potential interest in Candor Chasma, which we wanted to further investigate. Our secondary purpose was to collect additional geological samples from the stratification layers of the chasma.

Start time: 10:03 am

End time: 1:00 pm

Narrative: Today we decided to return to Candor Chasma, based on some interesting photos taken by EVA #6. We were looking for challenging terrain in which to test our rover, as well as sites of geological interest in the canyon. Driving along Cow Dung Road, we at first missed the Stream Bed Connector trail, as it is not marked and the recent weather has washed away part of the start of the trail. We realized once we hit Galileo Rd, turned our rovers around and headed back towards Stream Bed Connector, passing some very confused Martian tourists on the way. Confirming the start of Stream Bed Connector with our GPS and terrain map, we headed off towards the Chasma. EVA suit #7’s fan started to malfunction at the start of the EVA, delaying our start slightly, but was fixed with a quick power cycle. The walk along Stream Bed Connector took more than 20 minutes, as the sand was very moist and difficult to walk on, particularly with our heavy packs and equipment. We first took some time to explore along the lip of the Chasma, photographing the various rock strata and terrain types. It took some time to find an appropriate location to make our way down into the Chasma, and one of the EVA crew felt the terrain was too challenging for them to make their way down. It took another 20 minutes for the rest of the EVA team to find a suitable path down into the Chasma that all members could navigate. As we were running short on EVA time, the decision was made to focus primarily on collecting photos and locations of rover sites, rather than collection of geological samples. We left the sample equipment at the mouth of the Chasma to be picked up on our return, and explored roughly 1 km of the Chasma, taking photos of various rock overhangs, caves and terrain too challenging for a human in an EVA suit to explore. This will be perfect for rover exploration, provided the sand on the Chasma floor is not too moist! The weather today was extremely warm compared to previous EVAs (roughly 16 degrees) and we were extremely hot in our EVA suits, so stopped for water on the way back to the rovers, returning a little early due to the hot weather.

Destination: Walking along Stream Bed Connector and Candor Chasma.

Coordinates (use UTM NAD27 CONUS): 0520285/4251221

Participants: Jessica Todd (GreenHab Officer), Ghanim Aloitabi (Astronomer), Rawan Alshammari (HSO)

Road(s) and routes per MDRS Map: Drove along Entrance Road then north along Cow Dung Road (0110). Parked at entrance to Stream Bed Connector and walked along it to Candor Chasma

Mode of travel: Driving and Walking

Commander Report – February 08th

Maria Grulich


Crew 221, MDRS Sol 6

Crew 221 Commander Report 08FEB-2020

Author: Maria Grulich, Commander

Title: First problems to solve

Yesterday, during dinner, the crew 221 Earth2Mars realized that when the oven is on, there is a clear smell of propane more than usual. It was immediately reported to mission support during the comms window. At all times the crew was safe, as it was clear that when the oven is off, no propane gas could be smelled. This morning, after half the team went out for EVA, crew commander M.Grulich together with mission support started to investigate the source of the propane issue. It took a bit of time to read through the manual and check the oven, but after one hour the hypothesis was made that probably the oven was not configured to be used for propane gas. Per default the oven comes from the manufacturer to be used with natural gas instead of propane. In order to change this several grooves and an orifice spud on the stove and oven had to be replaced.

Mission support was very helpful, in supporting the theory and immediately organized support, such that in the afternoon, the oven was fixed with the help of the team. It was identified that the hypothesis was correct and after exchanging the parts, the gas issue seems to be solved. No propane smell can be detected now and also the gas detector was clear and the crew’s air quality detector improved significantly.

Besides this real issue, the team managed to do some science in the science dome, Greenhab and RAM.

The EVA team had some mishaps finding stream bed connector and they reported that the walk took more than 20 min as the soil was harder to walk on than expected. That resulted in a tired team that did not manage to explore the chasma as deep as they were planning on.

The team is in good shape and the cultural nights really add to the mission. We are making good progress on the projects; we are just not patient enough to see the results of the microbiology and ISRU plant experiments.

Commander Grulich

MDRS Station

Science Report – November 25th

Crew 221 Science Report 08-FEB-20
Crew Science Officer: Hemani Kalucha

1. Microbiology Research
Today, we tried to investigate the samples with bacteria growth under a microscope. One of the results was very exciting and is attached in the photo report.

2. ISRU project
Today, we tested the nutrient levels of the soil being used to grow black corn. We used a soil test kit, which involved putting in soil and water and a specific nutrient capsule (Phosphorous, Potassium, Nitrogen). The solution then changed to a specific shade after 10 minutes, and we could identify the state of soil healthy by looking at the coloured scale provided for each nutrient.

3. Mung Beans

We planted Mungbeans in 3 variations of soil, 100% martian soil, 50% martian soil and 50% earth soil, 70% earth soil and 30% martian soil. The soil used was from the EVA conducted at Condor Chasma yesterday. After planting, we inserted a conductivity probe and measured the salinity and water content of the soil with Arduino sensors. We need to convert the relative values to actual units.

4. Radio Communication

Crew Engineer constructed an antenna and got the radio to receive signals in the hab today! The antenna is 20 metres long, and it is in the RAM right now. It was listening at 40 m and 80 m wavelengths today, and managed to capture some interesting conversation all the way from Phoenix, Arizona: 465 miles away!

Astronomy Report – February 08th

Astronomy Report

Name: Ghanim Alotaibi

Crew: 221
Date: 08 Feb 2020

A Photometry measurement was submitted yesterday for V CAS (measurement performed 5 Feb). The V Mag was found to be 10.766 with an error of 0.0044.

The following table shows the new search pattern

(Re, Dec) (Re, Dec) (Re, Dec)
(15:00:00, +30:00:00) (15:03:00, +30:00:00) (15:06:00, +30:00:00)
(15:00:00, +30:45:00) (15:03:00, +30:45:00) (15:06:00, +30:45:00)
(15:00:00, +31:30:00) (15:03:00, +31:30:00) (15:06:00, +31:30:00)


Robotic Telescope Requested MDRS-14

Objects to be Imaged this Evening:

  1. V CAS
  2. Search pattern (submitted today): Only the first raw was submitted to familiarize the crew astronomer with the analysis procedure using Astroimage J.
  3. V0960 Mon (Submitted today).

Images submitted with this report:

  1. The attached image is the uncalibrated image taken by the MDRS-14 as per the request for Skynet mentioned in astronomy report 3 Feb 2020.
  2. The attached image is the uncalibrated image for V0960 MON.

Problems Encountered: The data pic for V0960 MON contains many stars with doughnut shape.

Sol Summary – February 08th

Crew 221 Sol Summary Report 08-Feb-2020


Summary Title: Well done…

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Update since last Sol Summary: Crew 221 played the rover game last night from the Planetary Volatiles Lab at York University, from Crew Journalist H. Kalucha before heading to an early night’s sleep.

Today, half the crew ( J.Todd, R.Alshammari, G. Alotaibi) went on an EVA towards Candor Chasma ,walking along Stream Bed Connector. The weather was beautiful and sunny and actually, the team reported that they were a bit too warm under the space suits. They did not achieve all their objectives but identified places to use the rover on future EVAs.

The other half of the team was staying in the hab. L. Monge and H.Kalucha were working in the science dome on soil health experiments, investigating the quality of the Martian soil being used to grow black corn. They attempted to determine if the black corn and mung beans have a chance to survive.

Commander M.Grulich was working on the oven issue, reading through the manuals and communicating with mission support to find a solution to our high propane measurements in the upper deck. The hypothesis was that the oven was not configured to be used with propane gas as the gas pressure regulator showed no change in configuration. Mission support was very helpful, checking information about the oven speedily, and eventually came to the same hypothesis.

In the afternoon, Crew 221 got external support and together with Dan, configured the oven to be used for propane. Since then, the flames of the stove are way smaller and the oven does not trigger the gas detector anymore. There are also no visible flames in the oven anymore, so the issue is solved!

This afternoon, the team worked on planting mung beans in the Greenhab that hopefully sprout soon. The crew already set a jar of mung beans aside that are sprouting since today to be hopefully ready to eat within the next couple of days.

Crew Engineer worked on his antenna project and managed to receive some signals from Arizona. Last but not least the microbiology experiments seemed to show first signs of growth on the soil samples. Tonight we will have Costa Rican pinto with a mocktail and the Costa Rican Cultural night, which we are all very excited about!

Look Ahead Plan: Sports Session, Meditation Session, Cultural night Costa Rica, Preparation rover and further microbiology samples and ISRU experiments

Anomalies in work: None

Update mouse in hab: still no joy in catching it.

Weather: sunny no winds in the morning, windy in the afternoon, no precipitation

Crew Physical Status: Good

EVA: 10:00an to 13:00pm

Reports to be filed:

Sol Summary

Commander Report

Photo Report

EVA Report

Operations Report

GreenHab Report

Science Report

Astronomy Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site